OpenStudy (anonymous):

find the number of PERMUTATIONS that can be formed from the letters of the word POPULAR and how many have two P's separated?

OpenStudy (anonymous):

if you had 7 different letters it would be \(7!\) by the counting principle. since you cannot tell the "P"s apart, it is \(\frac{7!}{2}\)

OpenStudy (anonymous):

well that gives the correct answer! thanks!!!

OpenStudy (anonymous):

good. now for the second part, i think we have to find the number of permutations where they ARE together and subtract that from the previous answer. does that seem right?

OpenStudy (anonymous):

yessss! plus, stationary salesman sell 1 pencil case, 1 eraser, 2 pens, 3 pencils,, howmany different sequences can these be sold?

OpenStudy (anonymous):

we can consider the 2 "P' s as one letter to compute that number. then it is like having 6 letters, and the number of ways to permute 6 letters is \(6!\)

OpenStudy (anonymous):

thank you, im really havin problem on this!!

OpenStudy (anonymous):

then your answer to part B should be \[\frac{7!}{2}-6!\] if i am not mistaken

OpenStudy (anonymous):

i get 1,800 http://www.wolframalpha.com/input/?i=7!%2F2-6! that is for part B of the first question now i will look at the second one

OpenStudy (anonymous):

i am not sure i understand this question stationary salesman sell 1 pencil case, 1 eraser, 2 pens, 3 pencils, how many different sequences can these be sold?

OpenStudy (anonymous):

if it means he can sell them in any order, but you cannot tell the pens apart, and cannot tell the pencils apart, then it is \[\frac{6!}{2!3!}\]

OpenStudy (anonymous):

does that look right?

OpenStudy (anonymous):

i am not sure if you can check your answers on line, but i interpret the question correctly it should be 60

OpenStudy (anonymous):

the answer is 420 in my scheme answer :/

OpenStudy (anonymous):

then i did not interpret the question correctly i wonder what it is actually asking

OpenStudy (anonymous):

oh i guess it would help if i could count! you have 7 items, not 6, so it is \[\frac{7!}{2!3!}=420\]

OpenStudy (anonymous):

i miscounted the number of items, sorry

OpenStudy (anonymous):

oh yes i got it , it is 7!\[\frac{ 7! }{ 2!3! }\]

OpenStudy (anonymous):

more importantly, is it clear how this works?

OpenStudy (anonymous):

yess.,, thank you :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

anyway for the first question, for the POPULAR, how many ways if 2 Ps are separated? because the answer on the top of the tread was for the "no of permutations can be formed from POPULAR"

OpenStudy (anonymous):

there are two parts to this question, right?

OpenStudy (anonymous):

1) number of permutations with no restriction we got as \(\frac{7!}{2!}=2520\)

OpenStudy (anonymous):

actually there is 4 parts of it, 1)no of permutations can be formed from POPULAR words 2) howmany ways begin and end with p 3)how many have 2P's separated 4)how many of them have vowels together

OpenStudy (anonymous):

and how is the working for no 3??

OpenStudy (anonymous):

for 3) it would be the number of ways with no restrictions, minus the number of ways with the two P together. that gives the total number of ways with the two separated. i answered that but will be happy to write it again

OpenStudy (anonymous):

we already computed the number of ways with no restrictions, it is \[\frac{7!}{2!}\] if the two P are together you can consider them as on letter "(PP)" and then it is like having 6 letters the number of ways to permute 6 letters is \(6!\) your answer is therefore \[\frac{7!}{2!}-6!\]

OpenStudy (anonymous):

yes!i overlooked, im sorry and thank you again,:)

OpenStudy (anonymous):

you good with the last one too?

OpenStudy (anonymous):

yup, i managed to answer that. hehe

OpenStudy (anonymous):

oh i thought that might be the trickiest don't you find some of these annoying? i do

OpenStudy (anonymous):

yeah, so annoying. i always that this topic is for geniuses! how you managed to get the working correctly? besides the fact that you are also a genius? is there any tricks or way to study them?sorry for the long post!